13 years ago, Shabbir's mother was 7 times as old as he was. She is now 48 years old. How many years old is Shabbir now?
Found 2 solutions by addingup, Alan3354:
Answer by addingup(3677)

Shabbir: S
Mother : M
M is now 48
---------------------
M-13 = 7(S-13)
48-13 = 7S-91
35 = 7S-91 flip the equation so we have the variable S on the left
7S-91 = 35
7S = 126
S = 18 So, Shabbir is presently 18
----------------------------------------
Check:
M-13 = 7(S-13)
48-13 = 7(18-13)
35 = 35 Correct.

Answer by Alan3354(69443)

13 years ago, Shabbir's mother was 7 times as old as he was. She is now 48 years old. How many years old is Shabbir now?
-----
13 years ago M was 35.
--> Shabbir was 5 13 yrs ago.
Now he's 18